Find the volume of a right circular cylinder with volume 1000 cm^3 and surface area 600 cm^2.
Answer by Alan3354(69443)

Find the volume of a right circular cylinder with volume 1000 cm^3 and surface area 600 cm^2.
------------
Volume = 1000 cubic cm.
Assuming there's no wall thickness, and that 600 cm^2 is the lateral area, not including the ends:
--------
Vol = pi*r^2*h = 1000
Area = 2pi*r*h = 600
----
h = 1000/(pi*r^2)
h = 600/2pi*r --> 300/(pi*r)
--------------
1000/(pi*r^2) = 300/(pi*r)
10/r^2 = 3/r
10 = 3r
r = 10/3 cm
------------------
h = 300/(pi*10/3)
h = 90/pi cm